OpenStudy (anonymous):

Write the equation of a circle with a center at (-2, -5) and a radius of 6.5.

OpenStudy (anonymous):

would it be (x + 2)2 + (y + 5)2 = 6.5 or (x + 2)2 + (y + 5)2 = 42.25

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

it would be ....= 6.5^2 = 42.25 :)